The measure of one angle is 47 degrees less that its supplement. What is the measure of each angle?
Answer by HyperBrain(694)

Let x=angle
Then 180-x=supplement of the angle
---
We have the equation:
x=(180-x)-47
x=133-x
2x=133
x=66.5 degrees